OpenStudy (diamond77108):

will medal & fan 2/9v+4−2/3v=5

OpenStudy (shadowlegendx):

\[\frac{ 2 }{ 9 }v + 4 - \frac{ 2 }{ 3 }v = 5\]

OpenStudy (shadowlegendx):

Do you know what the first step would be?

OpenStudy (diamond77108):

combine like terms

OpenStudy (shadowlegendx):

Correct, so how would we do that with these fractions?

OpenStudy (diamond77108):

\[\frac{ 2 }{ 9 }-\frac{ 2 }{ 3 }\]

OpenStudy (shadowlegendx):

Yes, but in order to add/subtract with fractions, what is required?

OpenStudy (diamond77108):

least common denominators which is 9

OpenStudy (shadowlegendx):

Yes, so how would we get 2/3 to fulfill that requirement?

OpenStudy (diamond77108):

\[\frac{ 6 }{ 9 }\]

OpenStudy (shadowlegendx):

Remember, \[-\frac{ 2 }{ 3 } \times \frac{ 3 }{ 3 } = - \frac{ 6 }{ 9 }\]

OpenStudy (shadowlegendx):

Negative

OpenStudy (shadowlegendx):

So we have \[\frac{ 2 }{ 9 }v + 4 - \frac{ 6 }{ 9 }v = 5\]

OpenStudy (shadowlegendx):

What now?

OpenStudy (diamond77108):

subtract the two fractions

OpenStudy (shadowlegendx):

What do you get?

OpenStudy (diamond77108):

\[\frac{ -8 }{ 9? }\]

OpenStudy (diamond77108):

positive 8

OpenStudy (shadowlegendx):

\[\frac{ 2 }{ 9 }v - \frac{ 6 }{ 9 }v\]

OpenStudy (diamond77108):

\[\frac{ -4 }{ 9 }\]

OpenStudy (shadowlegendx):

Yep \[-\frac{ 4 }{ 9 }v\] So we have, \[-\frac{ 4 }{ 9 }v + 4 = 5\]

OpenStudy (diamond77108):

so now we subtrac the 4 on both sides

OpenStudy (shadowlegendx):

Yes, We get.. \[-\frac{ 4 }{ 9 }v = 1\]

OpenStudy (diamond77108):

v=-4/9

OpenStudy (shadowlegendx):

Wait...

OpenStudy (shadowlegendx):

You multiply both sides by -9, you get 4v = -9, then divide by 4

OpenStudy (shadowlegendx):

\[v = -\frac{ 9 }{ 4 }\]

OpenStudy (diamond77108):

2 1/4

OpenStudy (shadowlegendx):

Keep it in the earlier form unless your teacher wants it in that form

OpenStudy (shadowlegendx):

Mixed fractions are kind of informal and used usually in just elementary school

OpenStudy (diamond77108):

its neg rigtht

OpenStudy (shadowlegendx):

Yes

OpenStudy (diamond77108):

thx

OpenStudy (shadowlegendx):

No problem